OpenStudy (anonymous):

- about star-twins what you know ? - bc. a star has five branch so than we use this 5 to generate star-twins - for example : 5-7 so 5+7=12 and now 12 +/- 5 = 7 - 17 what are twins too bc we know that 7 is twin with 5 and 17 is twin with 19 - so than this mean that the first STAR TWIN is 5-7 - so what will be the second star twin - can you getting it in this way ? @ganeshie8 ,@Nnesha ,@Mehek14,@TheSmartOne,@Kainui, @rebeccaxhawaii,@agent0smith,@zepdrix - opinions please - thank you in advance

rebeccaxhawaii (rebeccaxhawaii):

@Nnesha @Mehek14 @TheSmartOne @Kainui @agent0smith @zepdrix ^^^

OpenStudy (anonymous):

thank you

rebeccaxhawaii (rebeccaxhawaii):

never heard of a star twin

OpenStudy (anonymous):

@ganeshie8 after super twins here is something new

OpenStudy (anonymous):

so in this way we can generate the second star twin using 11 - 13 11 +13 = 24 +/ - 5 = 19 what is twin with 17 and 29 what is twin with 31 - yes ? so than this mean that 11 - 13 will be the second star twin

OpenStudy (anonymous):

@ganeshie8 please your opinion about ? ty

OpenStudy (anonymous):

so 17 - 19 work same too 17 + 19 = 36 +/- 5 = 31 what is twin with 29 and 41 what is twin with 43 how you like it ? - so than the 3rd star twin will be 17 - 19

OpenStudy (anonymous):

so what will be the next ?

OpenStudy (anonymous):

@jim_thompson5910 what is your opinion about this please ? thank you in advance

jimthompson5910 (jim_thompson5910):

@jhonyy9 did you mean twin primes? http://mathworld.wolfram.com/TwinPrimes.html

OpenStudy (anonymous):

yes but in this above wrote case star twin primes - generated in above wrote way

OpenStudy (anonymous):

@mathmale your opinion about this ? thank you

jimthompson5910 (jim_thompson5910):

sorry I've never heard of star primes or star twin primes

TheSmartOne (thesmartone):

I'm still trying to understand what a star twin is. (Can't find anything about it.) But it seems like to me that you're adding prime numbers and then subtracting 5 from it and then finding a twin prime 17 and 29 works 17 + 29 = 46 46 - 5 = 41 and 41 and 43 are twin prime numbers :P

OpenStudy (anonymous):

@TheSmartOne thank you that you try understanding the star twins - so how you see the name ,,star bc. how i have wrote on the posted text firstly bc a star has 5 branch so from this result that than you have a twin same with 5-7 and assume so result 12 so 12 +/- 5 generate 5 and 17 what mean that resulted two twins bc 5 is twin with 7 and 17 is twin with 19 so this mean that the twin 5-7 is star twin - this is same with ..super twins" than you have readed this my posted twin - generator how we can getting this super twins - so than you like i can tag you there for you can looking this too - wait your answer - jhonyy9

OpenStudy (anonymous):

@Kainui your opinion please about this ? thank you in advance

TheSmartOne (thesmartone):

sure, you can tag me

OpenStudy (anonymous):

ok - moment please

OpenStudy (anonymous):

@Kainui your opinion please about this - hope you remember what i've posted about super twins - so than what you think about star twins now ?

OpenStudy (kainui):

I see, it looks like you're connecting twin primes together. So twin primes are always of this form: \[6n\pm1\] for instance for n=3, \(17=6*3-1\) and \(19=6*3+1\) So what you're doing is taking your twin primes: \[(6n-1)+(6n+1)=12n\] then you're adding each of these to it: \[12n+6n-1 = 18n-1=6(3n)-1\]\[12n+6n+1 = 18n+1=6(3n)+1\] So if \(6n\pm 1\) are twin primes and \(6(3n)\pm 1\) are twin primes, then they're star twins. n=1 gives you \(6*1 \pm 1\) 5 and 7 and \(6*3*1 \pm 1\) gives you 17 and 19.

OpenStudy (anonymous):

ok and thank you @Kainui so after this generalized form do you can tel me what will be the next after this above wrote 3rd star twin so what are 17 and 19 - what will be the 4th star twin ?

ganeshie8 (ganeshie8):

Basically star twins are the twin primes such that 2p-3 and 2p+7 are primes. Super twins are the twin primes such that 2p+1 and 2p+3 are primes. Both these sets of primes seem to be very rare

ganeshie8 (ganeshie8):

71, 73 71 + 73 = 144 144 - 5 = 139 144 + 5 = 149

ganeshie8 (ganeshie8):

137, 139 149, 151 are pairs of twin primes

OpenStudy (anonymous):

ty @ganeshie8

OpenStudy (anonymous):

@ikram002p what is your opinion about these star twins ? and than you are here i like ask you that you have readed my before posted question,idea about super twins ?

OpenStudy (anonymous):

@heyyyyitsholly your opinion please about this ? thank you in advance